OpenStudy (anonymous):

To save material, Fred decides to build a horse corral using the barn for one side.He has 30m of fencing material and wants the corral to have an area of 100m^2. What are the dimensions of the corral.? Just need help with coming up with an equation

OpenStudy (anonymous):

the formula for the perimeter is P = 2l + 2w The formula for the area , A = lw so if the area needs to be 100 m^2 than 30m = 2l +2w and 100 = l w solve for l with the perimeter. so l = 15m -w substitute that back in the area formula. 100m^2 = (15 -w)(w) 100m = 15w -w^2 so now we have an equation w^2 -15w +100 = 0 can you solve this?

OpenStudy (anonymous):

ya just need to factor that out

OpenStudy (anonymous):

Good, do it than...

OpenStudy (anonymous):

14.11

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

What did you do?

OpenStudy (anonymous):

I used the quadratic formula

OpenStudy (anonymous):